# Breaking the Trend: How to Avoid Cherry-Picked Signals

### v10

@@ Appendix A Review of Literature of Solving Markowitz Optimization with Uncertainty of the Parameters
A complementary literature focuses on improving the estimation of the covariance matrix. Ledoit and Wolf (2003, 2004) developed optimal shrinkage estimators combining the sample covariance matrix with structured targets, leading to more stable and better-conditioned inputs for portfolio optimization. Jagannathan and Ma (2003) showed that imposing simple portfolio constraints (such as nonnegativity) implicitly regularizes the covariance matrix, mitigating estimation error effects. These approaches improve empirical robustness but, again, modify only the inputs of the optimization, not the analytical form of the optimal portfolio weights. Related structured-ensemble, predict-then-optimize methods steer diversification ex ante by tuning diversity in prediction and selection, using the ensemble’s combiner (equal-weight under MSE) as the target, while leaving the Markowitz closed form intact Rodriguez Dominguez at al (2025).
@@
In summary, the first term controls sensitivity to risk, the second captures expected performance under the nominal model, and the third explicitly regularizes against uncertainty in the mean estimates. Together, these terms yield a portfolio that trades off performance and robustness in a principled convex optimization framework. Segonne (2025) also uses $Q$ in a different way and introduces an anisotropy penalty into the mean-variance objective function to improve robustness. The term $\lambda\rho\,\|Q^{1/2}w\|_{2}$ is transformed into an anisotropy component through his Equation (48), which also helps interpret the solution of Benichou at al (2017) as a robust one. To our knowledge, this ”Robust Optimization and Worst-Case Formulations” is the only approach in the literature that explicitly accounts for the covariance matrix $Q$ of estimation errors, although the results differ from those of Benichou at al (2017) and Valeyre (2024).
@@
Grebenkov and Serror (2015) also optimises the portfolio by determining a matrix $A$ under the same proportionality assumption, but applies it specifically to the Grebenkov diffusive model and to exponential moving-average signals. Their results were derived independently from those of Brandt and Santa Clara (2006). Under suitable approximations, Valeyre (2024) showed that this matrix can be written as $C^{-1}C_{\xi}C^{-1}$, where $C_{\xi}$ is the covariance matrix of the estimation uncertainties $\xi$ of Eq. 1. By assuming a random diffusive model for $C_{\xi}$, Valeyre (2024) obtained the ARP portfolio, for which $C^{-1}C_{\xi}C^{-1}$ reduces to $C^{-1/2}$.
@@
Subsequently, Firoozye et al (2023); Kelly et al (2022) proposed simplified approximations from results of Brandt and Santa Clara (2006) leading to an easy-to-use ”generalised” Markowitz formula that properly incorporates $Q$ through their Eq. (6), namely $C^{-1}C_{rs}Q^{-1}$ (with our notations), instead of the $C^{-1}$ appearing in the classical Markowitz formula. Here, $C_{rs}$ denotes the ”cross” variance matrix between returns and signals.
@@
The expressions $C^{-1}C_{rs}Q^{-1}$ and $C^{-1}C_{\xi}C^{-1}$ become equivalent—or at least very close—under appropriate approximations when the signals are generated both by the Grebenkov diffusive process and by exponential moving averages. One may therefore say that Firoozye et al (2023); Kelly et al (2022) generalised the Markowitz formula, although this contribution has remained relatively unnoticed. Their work also generalises Grebenkov and Serror (2015), as it is not restricted to trend-following signals nor to the Grebenkov diffusion framework.
@@
All these unifying relationships between the different approaches of Grebenkov and Serror (2015); Kelly et al (2022); Firoozye et al (2023); Valeyre (2024); Benichou at al (2017) are clearly explained in Segonne (2025). Ultimately, it appears critical to model $C_{\xi}$ (or both $C_{rs}$ and $Q$) appropriately, since these matrices are largely unknown and cannot be measured with precision without causing overfiting to the portfolio. Modelling them through a random diffusive process naturally leads to the solution of Benichou at al (2017), ensuring that the resulting optimal portfolio is both robust and genuinely out-of-sample optimal. Empirical backtests, in the context of trend-folliwng, confirm that this approach delivers strong out-of-sample performance.
@@
Benichou at al (2017) and Valeyre (2024) therefore propose frameworks in which the optimal portfolio explicitly incorporates estimation uncertainty into its structure. These works derive solutions that differ fundamentally from the classical Markowitz prescription. In particular, under specific random matrix assumptions on the structure of the estimation-error correlation matrix, the optimal weights become proportional to $C^{-1/2}$ rather than $C^{-1}$, thus providing a substantial conceptual revision of the Markowitz approach. This line of research, where random matrix theory is obviously something crucial, remains rare in the literature but represents a promising direction for embedding measurement uncertainty at the core of portfolio theory.
@@ B.1 Acar (2003)
Acar (2003) uses a binary ”Long/Short” strategy (Eq. 17) in Section 8.5.1, ”Optimal long/short fund on the index itself XX” using his notations, where $F$ is the forecast used to predict the index $X$ and $H$ is the binary strategy. Acar (2003) presents his equation (Eq. 19) in Section 8.1, based on Acar (1998).
@@
In Eq. 19, we also need to derive $\rho_{xf}$, the correlation between the returns $X$ and the forecast $F$. This is not straightforward in the case of the simplest process which is used by Grebenkov and Serror (2014), to determine $\rho_{xf}$ in fonction of the paramters of the diffusive process and the parameter in the EMA used as the indicator. Julien Drouhet, my collegue in Machina Capital, derived the monstruous Eq. 39 and Eq. 40 in Sect B.2. We can not consider the formula of Eq. 19 as a ’elegant’ and not obvious closed solution as the $\rho_{xf}$ need to be determined through a long derivation.
@@ B.2 Closed formula derived from Acar (2003) and Grebenkov and Serror (2014) process
Given this dynamic, Grobenkov defines a momentum signal as an EMA of the returns as follows (Eq. (13), p.8):
@@ B.3 Acar’s formula for linear signals
We recall that $F$ is the forecast for the price of an instrument and $X$ is its future returns [^15] We keep the notations $\mu_{x},\mu_{f},\sigma_{x},\sigma_{f}$ and $\rho_{xf}$. The following result was proven by Firoozye and Koshiyama (2020).
@@
Ferreira et al. (2018), which cited Grebenkov and Serror (2014), used also a linear trading rule (Eq. 53) in their formula (3) from Section 2.2.2, ”Risk and Returns”.
@@
In Eq. 53, $m$ is a simple moving average (SMA) of returns rather than an exponential moving average (EMA). Ferreira et al. (2018) derived the Sharpe ratio over two pages without using the results of Acar (2003), obtaining an explicit formula in contrast to Acar (2003). In their formula (Eq. 54), which appears in Section 2.3, ”Limits and Interpretations”, the Sharpe ratio depends only on $\rho(t,t-i)$, the correlation coefficient of returns at times $t$ and tit-i. Ferreira et al. (2018) provided the formula closest to Grebenkov and Serror (2014). The formula is applicable to a more generalized diffusive process as in Grebenkov and Serror (2014), but it is less straightforward and not entirely consistent, since SMA differs from EMA. Ferreira et al. (2018) was published later and includes more parameters in the formula.
@@ B.5 Zakamulin and Giner (2022)
Zakamulin and Giner (2022) is also based on a simple moving average (SMA) of returns, Eq 55 (”2. Time Series Momentum Strategies”) but employs a binary strategy in eq 56. They derive the Sharpe ratio in Eq 59 using the same approach as Acar (2003), which requires the correlation between returns and the forecast ($\rho_{n}$ is ”the correlation coefficient between $r_{t}$ and the $MOM_{t-1}(n)$ ). This correlation is not straightforward to compute to claim explicit formula.
$\text{MOM}_{t-1}(n)=\sum_{i=1}^{n}X_{t-i}$ (55)
$R_{t}^{LS}=\begin{cases}r_{t},&\text{if }\text{MOM}_{t-1}(n)>0,\\[5.69054pt] 2r_{f,t}-r_{t},&\text{otherwise.}\end{cases}$ (56)

### v11

@@ Appendix A Review of Literature of Solving Markowitz Optimization with Uncertainty of the Parameters
A complementary literature focuses on improving the estimation of the covariance matrix. Ledoit and Wolf (2003, 2004) developed optimal shrinkage estimators combining the sample covariance matrix with structured targets, leading to more stable and better-conditioned inputs for portfolio optimization. Jagannathan and Ma (2003) showed that imposing simple portfolio constraints (such as nonnegativity) implicitly regularizes the covariance matrix, mitigating estimation error effects. These approaches improve empirical robustness but, again, modify only the inputs of the optimization, not the analytical form of the optimal portfolio weights. Related structured-ensemble, predict-then-optimize methods steer diversification ex ante by tuning diversity in prediction and selection, using the ensemble’s combiner (equal-weight under MSE) as the target, while leaving the Markowitz closed form intact Rodriguez Dominguez et al (2025).
@@
In summary, the first term controls sensitivity to risk, the second captures expected performance under the nominal model, and the third explicitly regularizes against uncertainty in the mean estimates. Together, these terms yield a portfolio that trades off performance and robustness in a principled convex optimization framework. Segonne (2025) also uses $Q$ in a different way and introduces an anisotropy penalty into the mean-variance objective function to improve robustness. The term $\lambda\rho\,\|Q^{1/2}w\|_{2}$ is transformed into an anisotropy component through his Eq. (48), which also helps interpret the solution of Benichou et al (2017) as a robust one. To our knowledge, this ”Robust Optimization and Worst-Case Formulations” is the only approach in the literature that explicitly accounts for the covariance matrix $Q$ of estimation errors, although the results differ from those of Benichou et al (2017) and Valeyre (2024).
@@
Grebenkov and Serror (2015) also optimises the portfolio by determining a matrix $A$ under the same proportionality assumption, but applies it specifically to the Grebenkov diffusive model and to exponential moving-average signals. Their results were derived independently from those of Brandt and Santa Clara (2006). Under suitable approximations, Valeyre (2024) showed that this matrix can be written as $A=C^{-1}C_{\xi}C^{-1}$, where $C_{\xi}$ is the covariance matrix of the estimation uncertainties $\xi$ of Eq. (1). By assuming a random diffusive model for $C_{\xi}$, Valeyre (2024) obtained the $\mathrm{ARP}$ portfolio, for which $A=C^{-1}C_{\xi}C^{-1}$ reduces to $A=C^{-1/2}$.
@@
Subsequently, Firoozye et al (2023); Kelly et al (2022) proposed simplified approximations from results of Brandt and Santa Clara (2006) leading to an easy-to-use ”generalised” Markowitz formula that properly incorporates $Q$ through their Eq. (6), namely $A=C^{-1}C_{rs}Q^{-1}$ (with our notations), instead of the $A=C^{-1}$ appearing in the classical Markowitz formula. Here, $C_{rs}$ denotes the ”cross” variance matrix between returns and signals.
@@
The expressions $A=C^{-1}C_{rs}Q^{-1}$ and $A=C^{-1}C_{\xi}C^{-1}$ become equivalent—or at least very close—under appropriate approximations when the signals are generated both by the Grebenkov diffusive process and by exponential moving averages. One may therefore say that Firoozye et al (2023); Kelly et al (2022) generalised the Markowitz formula, although this contribution has remained relatively unnoticed. Their work also generalises Grebenkov and Serror (2015), as it is not restricted to trend-following signals nor to the Grebenkov diffusion framework.
@@
All these unifying relationships between the different approaches of Grebenkov and Serror (2015); Kelly et al (2022); Firoozye et al (2023); Valeyre (2024); Benichou et al (2017) are clearly explained in Segonne (2025). Ultimately, it appears critical to model $C_{\xi}$ (or both $C_{rs}$ and $Q$) appropriately, since these matrices are largely unknown and cannot be measured with precision without causing overfiting to the portfolio. Modelling them through a random diffusive process naturally leads to the solution of Benichou et al (2017), ensuring that the resulting optimal portfolio is both robust and genuinely out-of-sample optimal. Empirical backtests, in the context of trend-following, confirm that this approach delivers strong out-of-sample performance.
@@
Benichou et al (2017) and Valeyre (2024) therefore propose frameworks in which the optimal portfolio explicitly incorporates estimation uncertainty into its structure. These works derive solutions that differ fundamentally from the classical Markowitz prescription. In particular, under specific random matrix assumptions on the structure of the estimation-error correlation matrix, the optimal weights become proportional to $A=C^{-1/2}$ rather than $A=C^{-1}$, thus providing a substantial conceptual revision of the Markowitz approach. This line of research, where random matrix theory is obviously something crucial, remains rare in the literature but represents a promising direction for embedding measurement uncertainty at the core of portfolio theory.
@@ B.1 Acar (2003)
Acar (2003) uses a binary ”Long/Short” strategy (Eq. (17) ) in Section 8.5.1, ”Optimal long/short fund on the index itself XX” using his notations, where $F$ is the forecast used to predict the index $X$ and $H$ is the binary strategy. Acar (2003) presents his equation (Eq. (19) ) in Section 8.1, based on Acar (1998).
@@
In Eq. (19), we also need to derive $\rho_{xf}$, the correlation between the returns $X$ and the forecast $F$. This is not straightforward in the case of the simplest process which is used by Grebenkov and Serror (2014), to determine $\rho_{xf}$ in fonction of the paramters of the diffusive process and the parameter in the EMA used as the indicator. Julien Drouhet, my collegue in Machina Capital, derived the monstruous Eq. (39) and Eq. (40) in Sect. B.2. We can not consider the formula of Eq. (19) as a ’elegant’ and not obvious closed solution as the $\rho_{xf}$ need to be determined through a long derivation.
@@ B.2 Closed formula derived from Acar (2003) and Grebenkov and Serror (2014) process
Given this dynamic, Grebenkov defines a momentum signal as an EMA of the returns as follows (Eq. (13), p.8):
@@ B.3 Acar’s formula for linear signals
We recall that $F$ is the forecast for the price of an instrument and $X$ is its future returns [^15] We keep the notations $\mu_{x},\mu_{f},\sigma_{x},\sigma_{f}$ and $\rho_{xf}$.. The following result was proven by Firoozye and Koshiyama (2020).
@@
Ferreira et al. (2018), which cited Grebenkov and Serror (2014), used also a linear trading rule (Eq. (53) ) in their formula (3) from Section 2.2.2, ”Risk and Returns”.
@@
In Eq. (53), $m$ is a simple moving average (SMA) of returns rather than an exponential moving average (EMA). Ferreira et al. (2018) derived the Sharpe ratio over two pages without using the results of Acar (2003), obtaining an explicit formula in contrast to Acar (2003). In their formula (Eq. (54) ), which appears in Section 2.3, ”Limits and Interpretations”, the Sharpe ratio depends only on $\rho(t,t-i)$, the correlation coefficient of returns at times $t$ and tit-i. Ferreira et al. (2018) provided the formula closest to Grebenkov and Serror (2014). The formula is applicable to a more generalized diffusive process as in Grebenkov and Serror (2014), but it is less straightforward and not entirely consistent, since SMA differs from EMA. Ferreira et al. (2018) was published later and includes more parameters in the formula.
@@ B.5 Zakamulin and Giner (2022)
Zakamulin and Giner (2022) is also based on a simple moving average (SMA) of returns, Eq. (55) (”2. Time Series Momentum Strategies”) but employs a binary strategy in Eq. (56). They derive the Sharpe ratio in Eq. (59) using the same approach as Acar (2003), which requires the correlation between returns and the forecast ($\rho_{n}$ is ”the correlation coefficient between $r_{t}$ and the $\mathrm{MOM}_{t-1}(n)$ ). This correlation is not straightforward to compute to claim explicit formula.
$\mathrm{MOM}_{t-1}(n)=\sum_{i=1}^{n}X_{t-i}$ (55)
$R_{t}^{LS}=\begin{cases}r_{t},&\text{if }\mathrm{MOM}_{t-1}(n)>0,\\[5.69054pt] 2r_{f,t}-r_{t},&\text{otherwise.}\end{cases}$ (56)